# Enhancing Logical Expressiveness in Graph Neural Networks via Path-Neighbor Aggregation

### v1

@@ Introduction
${\textbf{h}}_{v|u,q}^{(t+1)}=$ $\underset{w\in N_{r}(v),r\in R}{\text{UPDATE}}\bigg({\textbf{h}}_{v|u,q}^{(t)},\text{AGG}\Big(\{\{\text{MSG}_{r}({\textbf{h}}_{w|u,q}^{(t)},{\textbf{z}}_{\textbf{q}})\}\}\Big)\bigg)$
@@
The logical expressivity of a GNN refers to the set of logical formulas it can learn (understanding). C-GNN demonstrates superior logical expressive power compared to relational GNN (R-GNN), leading to enhanced performance (ATheory; understanding). We use the counting extension of modal logic (CML), which extends the counting expressiveness of graded modal logic, to describe FOL properties in KG. CML consists of propositional constants (always true $\top$, always false $\bot$), unary logical predicates $P_{i}(x)$, and recursive rules. Given FOL $\varphi(x,y)$, $\psi(x,y)$, positive integer variables $N\geq 1$ and $r\in R$, according to the recursive rules, $\neg\varphi(x,y)$, $\varphi(x,y)\wedge\psi(x,y)$, and $\exists^{N\geq n}z(\varphi(x,z)\wedge r(z,y))$ are also in CML. Based on the CML in KG, the logical rule expressive power of C-GNN can be expressed as follows:
@@
When an additional variable $z_{3}$ is introduced into T to break the symmetry, the two paths no longer share exactly the same intermediate nodes. We refer to this modified structure as U, as shown in Figure 3(d) and the corresponding rule formula is given as follows:
@@
(Adding Constants Restricts Coverage) If we start from the original rules $\varphi(x,y,\cdots)$ and replace at least one of them with a constant c to obtain new rules $\varphi{{}^{\prime}}(x,\textbf{c},\cdots)$, the coverage of the new rules is less than that of the original rules. Formally expressed as:
$Cov\Bigr(\varphi{{}^{\prime}}(x,\textbf{c},\cdots)\Bigr)\subseteq Cov\Bigr(\varphi(x,y,\cdots)\Bigr).$ (8)
@@
$\textbf{h}_{v|u,q}=\text{MLP}_{11}(\textbf{h}_{11})\odot\text{MLP}_{12}(\textbf{h}_{12})\odot\text{MLP}_{21}(\textbf{h}_{21})\odot\textbf{h}_{v|u,q}^{(L)}$ (11)
@@
To examine the effect of neighborhood hops on PN-GNN’s performance, we conducted ablation studies on $T$, $U$, FB15K237 v1, and WN18RR v1, comparing $\text{PN-GNN}_{11}$ (using only 1-hop) and $\text{PN-GNN}_{12-21}$ (using only 2-hop). Results are shown in Table 4. C-GNN is sufficient for the $T$ dataset, with all models achieving 100% accuracy. For $U$, $\text{PN-GNN}_{12-21}$ outperforms NBFNet by 14.8%, consistent with Lemma 8, as distinguishing the logical rules in $U$ mainly relies on 3-hop neighborhoods. On FB15K237 v1, PN-GNN and its variants all outperform NBFNet, with the best improvement reaching 5%, demonstrating the benefits of improved logical expressiveness in complex real-world datasets. On WN18RR v1, $\text{PN-GNN}_{11}$ achieves the highest performance, outperforming NBFNet by 2.7%. This suggests that for datasets with fewer relations and predominantly short-path logical structures, leveraging only 1-hop neighborhoods provides more relevant information than considering longer paths, which may introduce noise.
Conclusion and Future work

### v2

@@ Introduction
${\textbf{h}}_{v|u,q}^{(t+1)}=$ $\underset{\begin{subarray}{c}w\in N_{r}(v)\\ r\in R\end{subarray}}{\text{UPDATE}}\bigg({\textbf{h}}_{v|u,q}^{(t)},\text{AGG}\Big(\{\{\text{MSG}_{r}({\textbf{h}}_{w|u,q}^{(t)},{\textbf{z}}_{\textbf{q}})\}\}\Big)\bigg)$
@@
The logical expressivity of a GNN refers to the set of logical formulas it can learn (understanding). C-GNN demonstrates superior logical expressive power compared to R-GNN, leading to enhanced performance (ATheory; understanding). We use the counting extension of modal logic (CML), which extends the counting expressiveness of graded modal logic, to describe FOL properties in KG. CML consists of propositional constants (always true $\top$, always false $\bot$), unary logical predicates $P_{i}(x)$, and recursive rules. Given FOL $\varphi(x,y)$, $\psi(x,y)$, positive integer variables $N\geq 1$ and $r\in R$, according to the recursive rules, $\neg\varphi(x,y)$, $\varphi(x,y)\wedge\psi(x,y)$, and $\exists^{N\geq n}z(\varphi(x,z)\wedge r(z,y))$ are also in CML. Based on the CML in KG, the logical rule expressive power of C-GNN can be expressed as follows:
@@
When the relations $r_{2}$ and $r_{4}$ are outgoing edges of the same node, we refer to this modified structure as U, as shown in Figure 3(d) and the corresponding rule formula is given as follows:
@@
(Adding Constants Restricts Coverage) If we start from the original rules $\varphi(x,y,\cdots)$ and replace at least one of them with a constant c to obtain new rules $\varphi{{}^{\prime}}(x,c,\cdots)$, the coverage of the new rules is less than that of the original rules. Formally expressed as:
$Cov\Bigr(\varphi{{}^{\prime}}(x,c,\cdots)\Bigr)\subseteq Cov\Bigr(\varphi(x,y,\cdots)\Bigr).$ (8)
@@
$\textbf{h}_{d}=\text{MLP}_{11}(\textbf{h}_{11})\odot\text{MLP}_{12}(\textbf{h}_{12})\odot\text{MLP}_{21}(\textbf{h}_{21})$ (11)
$\textbf{h}_{v|u,q}=\textbf{h}_{d}\odot\textbf{h}_{v|u,q}^{(L)}$
@@
To examine the effect of neighborhood hops on PN-GNN’s performance, we conducted ablation studies on $T$, $U$, FB15K237 v1, and WN18RR v1, comparing $\text{PN-GNN}_{11}$ (using only 1-hop) and $\text{PN-GNN}_{12-21}$ (using only 2-hop). Results are shown in Table 4. C-GNN is sufficient for the $T$ dataset, with all models achieving 100% accuracy. For $U$, $\text{PN-GNN}_{12-21}$ outperforms NBFNet by 15.8%, consistent with Lemma 8, as distinguishing the logical rules in $U$ mainly relies on 3-hop neighborhoods. On FB15K237 v1, PN-GNN and its variants all outperform NBFNet, with the best improvement reaching 5%, demonstrating the benefits of improved logical expressiveness in complex real-world datasets. On WN18RR v1, $\text{PN-GNN}_{11}$ achieves the highest performance, outperforming NBFNet by 2.7%. This suggests that for datasets with fewer relations and predominantly short-path logical structures, leveraging only 1-hop neighborhoods provides more relevant information than considering longer paths, which may introduce noise.
Conclusion